OpenStudy (anonymous):

2^(4/log_2(x)=1/256 solve for x

OpenStudy (anonymous):

\[2^{4/l \log_{2}x}=1/256\]

OpenStudy (campbell_st):

1/256 = 2^(-8) equating powers 2^(-8) = 2^2/log(x) Log(x) = 2^2/2^(-8) log(x) = 2^(10) then x = (2)^(2^10)

OpenStudy (anonymous):

which is?

OpenStudy (campbell_st):

2^1024

OpenStudy (anonymous):

it says to answer to 4 decimal places...